(x-2)/(x-3)=(x-7)/(x-9)
How do I solve this problem????

Answer by Nate(3500)

(x - 2)(x - 9) = (x - 7)(x - 3)
x^2 - 11x + 18 = x^2 - 10x + 21
-3 = x